HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

POPULATING THE EXOMOL DATABASE WITH LINESHAPE PARAMETERS: COLLISIONAL BROADENING AND SHIFTING OF

ACETYLENE

\underline{A. SOKOLOV} , S.N. YURCHENKO, J. TENNYSON, Department of Physics and

Astronomy, University College London, London WC1E 6BT, UK; R.R. GAMACHE, Department of Environmental, Earth, and Atmospheric Sciences,

University of Massachusetts Lowell, Lowell, MA 01854, USA

Collisional lineshape parameters are one of the indispensable data inputs necessary to produce absorption/emission cross-sections which are required in radiative transfer calculations and spectral retrievals. To be exhaustive, these parameters need to be calculated for a number of “radiating molecule – collisional perturber” pairs and at different temperatures. Of particular interest to a burgeoning field of exoplanetary science are the parameters at temperatures higher than those usually required for terrestrial studies, typically around 500–3000 K.

We discuss our recent progress in using a semiclassical method called Modified Complex Robert-Bonamy (MCRB) approach in calculating broadening and shifting coefficients of acetylene (C_{2}H_{2}) for several perturbers (H_{2}, He, N_{2} and C_{2}H_{2}) for a large range of temperatures. This report is a part of the ongoing work to populate the ExoMol database with lineshape parameters, including the line list for acetylene aCeTy containing over 3 billion transitions, with little or no line broadening information. Comparisons with a large pool of existing theoretical and experimental acetylene data are shown. In addition, discussed are the difficulties arising when modelling vibrational dependence of lineshape parameters.

p-number: p199  Submitted on Thu, 15 Jun 2023 19:13:13 +0200